# A note on the cohomology of moduli spaces of local shtukas

### v1

@@ 1. Introduction
The purpose of this note is to explicate a formula for the cohomology of moduli spaces of local shtukas that can be derived from the recent work of Fargues–Scholze [FS21], and to note some consequences. Let $p\neq\ell$ be distinct primes. The theory of local Shimura varieties, which began with examples such as the Lubin–Tate and Drinfeld towers, and continued with the moduli spaces of $p$-divisible groups of Rapoport–Zink [RZ96], has reached a stage of maturity exceeding that of the global theory of Shimura varieties with Scholze’s general definition of moduli spaces of local shtukas [SW20]. For any connected reductive group $G/\mathbb{Q}_{p}$, any element $b$ in Kottwitz’s set $B(G)$, and any finite collection $\mu_{\bullet}=(\mu_{i})_{i\in I}$ of conjugacy classes of cocharacters $\mu_{i}$ of $G$, Scholze defines a tower
@@
of diamonds, with inverse limit $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$, where $K\subseteq G(\mathbb{Q}_{p})$ runs through the compact open subgroups. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter, the $\operatorname{Sht}_{(G,b,\mu),K}$ are smooth rigid spaces referred to as local Shimura varieties. The space $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$ carries commuting actions of $G(\mathbb{Q}_{p})$ and $G_{b}(\mathbb{Q}_{p})$, where $G_{b}$ is the inner form of a Levi subgroup of the quasisplit form of $G$ canonically attached to the datum $(G,b)$. For any irreducible admissible $G_{b}(\mathbb{Q}_{p})$-representation $\rho$ over $\overline{\mathbb{Q}}_{\ell}$, one may define the “$\rho$-isotypic part of the $\overline{\mathbb{Q}}_{\ell}$-intersection cohomology of $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$”, which we will denote by
@@
We note that this definition naturally incorporates a shift making 00 the ‘middle degree’. It carries commuting actions of $G(\mathbb{Q}_{p})$ and $W_{E_{\bullet}}:=\prod_{i\in I}W_{E_{i}}$ (where $W_{E_{i}}$ is the Weil group of the reflex field $E_{i}$ of $\mu_{i}$) and is a bounded complex of finite length admissible $G(\mathbb{Q}_{p})$-representations. Roughly speaking, the local Langlands conjecture associates an $L$-parameter $\phi=\phi_{\rho}:W_{\mathbb{Q}_{p}}\to(\widehat{G}\rtimes W_{\mathbb{Q}_{p}})(\overline{\mathbb{Q}}_{\ell})$ with $\rho$, as well as an $L$-packet $\Pi_{\phi}(G)$ of irreducible admissible representations of $G(\mathbb{Q}_{p})$. Furthermore, there should be a relation between $\operatorname{Irr}(S_{\phi},\chi_{b})$ and $\Pi_{\phi}(G)$. Here $S_{\phi}$ is the centralizer in $\widehat{G}$ of the image of $\phi$, and $\operatorname{Irr}(S_{\phi},\chi_{b})$ denotes the set of irreducible representations of $S_{\phi}$ on which $Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}\subseteq S_{\phi}$ acts by a certain character $\chi_{b}$ determined by $b$. Let $V$ be the dual of the irreducible representation of $\prod_{i\in I}\widehat{G}\rtimes W_{E_{i}}$ with extreme weight $\mu_{\bullet}$. The underlying vector space of $V$ also carries an action of $W_{E_{\bullet}}$ coming from $\phi$, and this action makes it into a $S_{\phi}\times W_{E_{\bullet}}$-representation that we will denote by $V_{\phi}$. Define
@@
in the Grothendieck group of $G(\mathbb{Q}_{p})\times W_{E_{\bullet}}$-representations. The following is then the natural generalization of the Kottwitz conjecture (see e.g. [RV14, §7.1]), together with a folklore vanishing conjecture.
@@
Assume that $b$ is basic and $\phi=\phi_{\rho}$ is elliptic, i.e. $\phi$ is semisimple and $S_{\phi}/Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}$ is finite. In this case the local Langlands correspondence predicts a surjection $\operatorname{Irr}(S_{\phi},\chi_{b})\to\Pi_{\phi}(G)$, which we denote by $\delta\mapsto\pi_{\delta}$.
@@
    (Kottwitz conjecture) We have $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})}\pi_{\delta}\boxtimes\operatorname{Hom}_{S_{\phi}}(\delta,V_{\phi})$.
@@
$R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}C_{\delta}\boxtimes\operatorname{Hom}_{S_{\varphi}}(\delta,V_{\varphi}),$
@@
The precise version, which notably includes a formula for each $C_{\delta}$ in terms of $\rho$ and $\delta$, is given in Theorem 3.5 and Corollary 3.8 (and works in positive characteristic as well). A parameter $\varphi$ is called generous if it is semisimple, if no other $L$-parameter has semisimplification $\varphi$, and if $\varphi$ satisfies an additional technical moduli-theoretic condition (see Definition 3.1). Generous parameters are generic on the coarse moduli space and includes the elliptic parameters. Thus, under the expectation that $\varphi=\phi$, the action of $W_{E_{\bullet}}$ on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is as predicted from the Kottwitz conjecture. We note that our formula may be seen as a (more general) local analogue of [LZ19, Prop. 1.2].
The proof of Theorem 1.2 is given in sections 2 and 3. The main idea is to apply the machinery of higher algebra to the spectral action of Fargues–Scholze, to obtain a version of the spectral action which only sees one $L$-parameter at a time. The general version of this idea is described in §2, and in §3 we use to derive Theorem 1.2. The key point that we wish to make, and which is needed to carry out the proof, is that even though the machinery that we use (monoidal $\infty$-categories and their modules) is highly abstract, it allows you to make computations. This would fail if we tried to work with triangulated categories instead of their $\infty$-categorical enhancements.
Section 4 then gives some applications of Theorem 1.2 to both parts of Conjecture 1.1; we highlighting one such application. When disregarding the $W_{E_{\bullet}}$-action, Conjecture 1.1 was recently proven in [HKW22] under the assumption of a precise form of the local Langlands correspondence. Combining this with Theorem 1.2, one gets the following result.
@@
Assume the refined local Langlands correspondence [Kal16, Conjecture G], and let $\phi$ be the $L$-parameter attached to $\rho$. Assume further that the Fargues–Scholze parameter $\varphi$ attached to $\rho$ is elliptic, that all $\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})$ are one-dimensional, and that all representations in $\Pi_{\phi}(G)$ are supercuspidal. Then there exists a surjection $\delta\mapsto\pi_{\delta}^{\prime}$ from $\operatorname{Irr}(S_{\varphi},\chi_{b})$ to $\Pi_{\phi}(G)$ such that
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\pi_{\delta}^{\prime}\boxtimes V_{\varphi,i}$
@@
In particular, this is the Kottwitz conjecture, up to the comparison of the local Langlands correspondence and the Fargues–Scholze construction.
@@ Acknowledgements
C.J. was supported by Vetenskapsrådet Grant 2020-05016, Geometric structures in the p-adic Langlands program. D.H. was supported by a startup grant through the National University of Singapore.
@@ 2. Spectral action with supports
The main player on the geometric side of the geometrization of the local Langlands correspondence is the stable $\infty$-category $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\overline{\mathbb{Q}}_{\ell})$ defined in [FS21, §VII.7] and its counterparts $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{Q}}_{\ell})$ on the strata $\mathrm{Bun}_{G}^{b}$, which are equivalent to the derived categories $\mathcal{D}(G_{b}(E),\overline{\mathbb{Q}}_{\ell})$ of smooth representations of $G_{b}(E)$ over $\overline{\mathbb{Q}}_{\ell}$.
@@
which are functorial as $I$ varies.
On the spectral side, we have the stack of $L$-parameters $Z^{1}(W_{E},\widehat{G})/\widehat{G}$ over $\overline{\mathbb{Q}}_{\ell}$, defined in [FS21, §VIII]. Since it will figure extensively in our discussion and formulas, we simplify the notation by defining
@@
We also let $X^{\Box}:=Z^{1}(W_{E},\widehat{G})$ be the representation variety and let $X:=Z^{1}(W_{E},\widehat{G})\sslash\widehat{G}$ be its GIT quotient (the character variety), which is a good moduli space for $\mathfrak{X}$. The spectral Bernstein center $\mathcal{Z}^{\mathrm{spec}}(G):=\mathcal{O}(X^{\Box})^{\widehat{G}}$ is then the ring of global functions on $\mathfrak{X}$, and we let $\operatorname{Perf}(\mathfrak{X})$ denote the stable $\infty$-category of perfect complexes on $\mathfrak{X}$. By [FS21, Thm. X.0.1], the action of the Hecke operators on $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ is equivalent to an exact $\overline{\mathbb{Q}}_{\ell}$-linear monoidal functor
@@
called the spectral action. Here and below $\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}$ denotes the $\infty$-category of exact $\overline{\mathbb{Q}}_{\ell}$-linear endofunctors, and $\operatorname{End}^{L}_{\overline{\mathbb{Q}}_{\ell}}\subseteq\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}$ denotes the full subcategory of colimit-preserving endofunctors.
@@
$\operatorname{Rep}((\widehat{G}\rtimes Q)^{I})\to\operatorname{Perf}(\mathfrak{X})^{BW_{E}^{I}}\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}))^{\omega})^{BW_{E}^{I}},$
@@
to the geometric Bernstein center $\mathcal{Z}^{\mathrm{geom}}(G)$ of $G$, landing inside the subring $\mathcal{Z}^{\mathrm{geom}}_{\mathrm{Hecke}}(G)$ of endomorphisms equivariant for the Hecke action [FS21, Def. IX.0.2]. For any object $A\in\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ we have a ring homomorphism $F_{A}:\mathcal{Z}^{\mathrm{spec}}(G)\to\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)$. When $A$ is Schur (i.e. $\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)=\overline{\mathbb{Q}}_{\ell}$), the kernel of $F_{A}$ corresponds to a $\overline{\mathbb{Q}}_{\ell}$-point of $X$, which we refer to as the Fargues–Scholze parameter of $A$. By [Zou22, Theorem 5.2.1], this construction is compatible with that given in [FS21]. For $b\in B(G)$ and $\rho$ an irreducible admissible representation of $G_{b}(E)$, we will write “the Fargues–Scholze parameter of $\rho$” to mean the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ [^2] We note that using $i_{\natural}^{b}\rho$ instead of $i^{b}_{\ast}\rho$, as is done in [FS21, IX.7], produces the same result (e.g. by the argument in the proof of Proposition 2.6). Moreover, the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ is the composite of the “correct” Fargues–Scholze parameter of $\rho$ (defined using $i^{1}_{!}\rho$ in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G_{b}})$ as in [FS21, Def. IX.7.1]) and the twisted $L$-embedding $\widehat{G}_{b}\rtimes Q\to\widehat{G}\rtimes Q$, by [FS21, Thm IX.7.2]..
Now consider a derived stack $\mathfrak{Y}$ with a map $f:\mathfrak{Y}\to\mathfrak{X}$, which induces a monoidal functor $f^{\ast}:\operatorname{QCoh}(\mathfrak{X})\to\operatorname{QCoh}(\mathfrak{Y})$. Using the spectral action and $f^{\ast}$, we can then form the tensor product
$\operatorname{QCoh}(\mathfrak{Y})\otimes_{\operatorname{QCoh}(\mathfrak{X})}D_{\mathrm{lis}}(\mathrm{Bun}_{G})$
@@
To keep the definitions of this paper close to its theorems, we make our definitions in a restricted setting, where things are more concrete. Consider the character variety $X$ from above. Let $Y$ be a closed subscheme of $X$, and consider its derived pullback $\mathfrak{Y}$ to $\mathfrak{X}$. The structure sheaf $\mathcal{O}_{\mathfrak{Y}}$ is a commutative algebra object of $\operatorname{QCoh}(\mathfrak{X})$. Similarly, the structure sheaf $\mathcal{O}_{Y}$ of $Y$ is a commutative algebra object of $\operatorname{QCoh}(X)$, and pullback gives a natural monoidal functor $\operatorname{QCoh}(X)\to\operatorname{QCoh}(\mathfrak{X})$ sending $\mathcal{O}_{Y}$ to $\mathcal{O}_{\mathfrak{Y}}$. In particular, we may and will also view $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ as module for $\operatorname{QCoh}(X)$.
@@
We start with part (1). For any $k$-algebras $R$ and $S$, the $\infty$-category of colimit-preserving functors $\mathcal{D}(R)\to\mathcal{D}(S)$ is equivalent to $\mathcal{D}(S\otimes_{k}R^{op})$, or equivalently to the derived $\infty$-category $\mathcal{D}(S,R)$ of $(S,R)$-bimodules. Here $X\in\mathcal{D}(S,R)$ corresponds to the functor $M\mapsto X\otimes_{R}^{L}M$, and for a given $F:\mathcal{D}(R)\to\mathcal{D}(S)$ the corresponding $X$ is $F(R)$, with the left $R^{op}$-module structure coming from the map
@@
Since $A$ and $B$ are concentrated in degree 00, this map factors through the center $Z(B)=\pi_{0}(\operatorname{End}_{\mathcal{D}(B,B)}(B))$, giving the desired map $f$. Translating back from $\mathcal{D}(B,B)$ to $\operatorname{End}_{k}^{L}(\mathcal{D}(B))$ then gives the desired formula.
For part (2), by shifting we may assume that $M$ has homology in degree 00, and so lies in the heart of $\mathcal{D}(B)$ with respect to the usual t-structure. The assumption then says that $M$ has a canonical $B/IB$-module structure in the usual sense. Restriction along the canonical map $A/I\otimes^{L}_{A}B\to B/IB$ then gives a canonical $A/I\otimes_{A}^{L}B$-module structure, as desired. ∎

### v2

@@ 1. Introduction
The purpose of this note is to explicate a formula for the cohomology of moduli spaces of local shtukas that can be derived from the recent work of Fargues–Scholze [FS21], and to note some consequences. Let $p\neq\ell$ be distinct primes. The theory of local Shimura varieties, which began with examples such as the Lubin–Tate and Drinfeld towers, and continued with the moduli spaces of $p$-divisible groups of Rapoport–Zink [RZ96], has reached a new level of generality with Scholze’s definition of moduli spaces of local shtukas [SW20]. For any connected reductive group $G/\mathbb{Q}_{p}$, any element $b$ in Kottwitz’s set $B(G)$, and any finite collection $\mu_{\bullet}=(\mu_{i})_{i\in I}$ of conjugacy classes of cocharacters $\mu_{i}$ of $G$, Scholze defines a tower
@@
of diamonds, with inverse limit $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$, where $K\subseteq G(\mathbb{Q}_{p})$ runs through the compact open subgroups. When $\mu_{\bullet}=\mu$ is a single minuscule cocharacter, the $\operatorname{Sht}_{(G,b,\mu),K}$ are smooth rigid spaces referred to as local Shimura varieties. The space $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$ carries commuting actions of $G(\mathbb{Q}_{p})$ and $G_{b}(\mathbb{Q}_{p})$, where $G_{b}$ is the inner form of a Levi subgroup of the quasisplit form of $G$ canonically attached to the datum $(G,b)$. For any irreducible admissible $G_{b}(\mathbb{Q}_{p})$-representation $\rho$ over $\overline{\mathbb{Q}}_{\ell}$, one may define the $\rho$-isotypic part of the $\overline{\mathbb{Q}}_{\ell}$-intersection cohomology of $\operatorname{Sht}_{(G,b,\mu_{\bullet})}$”, which we will denote by
@@
We note that this definition naturally incorporates a shift making 0 the ‘middle degree’. It carries commuting actions of $G(\mathbb{Q}_{p})$ and $W_{E_{\bullet}}:=\prod_{i\in I}W_{E_{i}}$ (where $W_{E_{i}}$ is the Weil group of the reflex field $E_{i}$ of $\mu_{i}$) and is a bounded complex of finite length admissible $G(\mathbb{Q}_{p})$-representations. Roughly speaking, the local Langlands conjecture associates an $L$-parameter $\phi=\phi_{\rho}:W_{\mathbb{Q}_{p}}\to(\widehat{G}\rtimes W_{\mathbb{Q}_{p}})(\overline{\mathbb{Q}}_{\ell})$ with $\rho$, as well as an $L$-packet $\Pi_{\phi}(G)$ of irreducible admissible representations of $G(\mathbb{Q}_{p})$. Furthermore, there should be a relation between $\operatorname{Irr}(S_{\phi},\chi_{b})$ and $\Pi_{\phi}(G)$. Here $S_{\phi}$ is the centralizer in $\widehat{G}$ of the image of $\phi$, and $\operatorname{Irr}(S_{\phi},\chi_{b})$ denotes the set of irreducible representations of $S_{\phi}$ on which $Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}\subseteq S_{\phi}$ acts by a certain character $\chi_{b}$ determined by $b$. Let $V$ be the dual of the irreducible representation of $\prod_{i\in I}\widehat{G}\rtimes W_{E_{i}}$ with extreme weight $\mu_{\bullet}$. The underlying vector space of $V$ also carries an action of $W_{E_{\bullet}}$ coming from $\phi$, and this action makes it into a $S_{\phi}\times W_{E_{\bullet}}$-representation that we will denote by $V_{\phi}$. Define
@@
in the Grothendieck group of $G(\mathbb{Q}_{p})\times W_{E_{\bullet}}$-representations. Write $\operatorname{Rep}(S_{\phi},\chi_{b})$ for the set of finite-dimensional representations of $S_{\phi}$ on which $Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}$ acts by $\chi_{b}$. The following is then the natural generalization of the Kottwitz conjecture (see e.g. [RV14, §7.1], [HKW22, Conjecture 1.0.1]), together with a folklore vanishing conjecture.
@@
Assume that $b$ is basic and $\phi=\phi_{\rho}$ is elliptic, i.e. $\phi$ is semisimple and $S_{\phi}/Z(\widehat{G})^{W_{\mathbb{Q}_{p}}}$ is finite. In this case a construction assuming the local Langlands correspondence gives a map $\Pi_{\phi}(G)\to\operatorname{Rep}(S_{\phi},\chi_{b})$, which we denote by $\pi\mapsto\delta_{\pi,\rho}$.
@@
    (Kottwitz conjecture) We have $\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\pi\in\Pi_{\phi}(G)}\pi\boxtimes\operatorname{Hom}_{S_{\phi}}(\delta_{\pi,\rho},V_{\phi})$.
@@
$R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\bigoplus_{\delta\in\operatorname{Irr}(S_{\phi},\chi_{b})}C_{\delta}\boxtimes\operatorname{Hom}_{S_{\varphi}}(\delta,V_{\varphi}),$
@@
The precise version, which notably includes a formula for each $C_{\delta}$ in terms of $\rho$ and $\delta$, is given in Theorem 3.8 and Corollary 3.11 (and works in positive characteristic as well). A parameter $\varphi$ is called generous if it is semisimple, if no other $L$-parameter has semisimplification $\varphi$, and if $\varphi$ satisfies an additional technical moduli-theoretic condition (see Definition 3.2). Generous parameters are generic on the coarse moduli space and includes the elliptic parameters. Thus, under the expectation that $\varphi=\phi$, the action of $W_{E_{\bullet}}$ on $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is essentially as predicted from the Kottwitz conjecture. We note that our formula may be seen as a (more general) local analogue of [LZ19, Prop. 1.2].
The proof of Theorem 1.2 is given in sections 2 and 3. The main idea is to apply the machinery of higher algebra to the spectral action of Fargues–Scholze, to obtain a version of the spectral action which only sees one $L$-parameter at a time. The general version of this idea is described in §2, and we believe that this should be a useful tool in the study of the spectral action in general. In §3 we use it to derive Theorem 1.2. The key point that we wish to make, and which is needed to carry out the proof, is that even though the machinery that we use (monoidal $\infty$-categories and their modules) is highly abstract, it allows you to make computations. This would fail if we tried to work with triangulated categories instead of their $\infty$-categorical enhancements.
Section 4 then gives some applications of Theorem 1.2 to both parts of Conjecture 1.1; we highlighting one such application here. When disregarding the $W_{E_{\bullet}}$-action, Conjecture 1.1 was recently proven in [HKW22] under the assumption of a precise form of the local Langlands correspondence. Combining this with Theorem 1.2, one gets the following result.
@@
Assume the refined local Langlands correspondence [Kal16, Conjecture G], and let $\phi$ be the $L$-parameter attached to $\rho$. Assume further that the Fargues–Scholze parameter $\varphi$ attached to $\rho$ is elliptic, that all $\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})$ are one-dimensional, and that all representations in $\Pi_{\phi}(G)$ are supercuspidal. Then there exists a surjection $\delta\mapsto\pi_{\delta}$ from $\operatorname{Irr}(S_{\varphi},\chi_{b})$ to $\Pi_{\phi}(G)$ such that
$\operatorname{Mant}_{G,b,\mu_{\bullet}}(\rho)=\sum_{\delta\in\operatorname{Irr}(S_{\varphi},\chi_{b})}\pi_{\delta}\boxtimes\operatorname{Hom}_{S_{\varphi}}(\delta,V_{\varphi})$
@@ Acknowledgements
We thank the anonymous referee for many comments which helped improve the paper. We also thank Pol van Hoften for helpful discussions on a previous version of this paper. C.J. was supported by Vetenskapsrådet Grant 2020-05016, Geometric structures in the p-adic Langlands program. D.H. was supported by a startup grant through the National University of Singapore.
@@ 2. Spectral action with supports
The main player on the geometric side of the geometrization of the local Langlands correspondence is the stable $\infty$-category $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\overline{\mathbb{Q}}_{\ell})$, defined in [FS21, §VII.7], and its counterparts $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G}^{b},\overline{\mathbb{Q}}_{\ell})$ on the strata $\mathrm{Bun}_{G}^{b}$, which are equivalent to the derived categories $\mathcal{D}(G_{b}(E),\overline{\mathbb{Q}}_{\ell})$ of smooth representations of $G_{b}(E)$ over $\overline{\mathbb{Q}}_{\ell}$.
@@
which are functorial as $I$ varies. Here and below, whenever $\mathcal{C}$ is a stable $\overline{\mathbb{Q}}_{\ell}$-linear $\infty$-category, we let $\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{C})$ denote the $\infty$-category of exact $\overline{\mathbb{Q}}_{\ell}$-linear endofunctors of $\mathcal{C}$. Moreover, $\operatorname{End}^{L}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{C})\subseteq\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{C})$ will denote the full subcategory of colimit-preserving endofunctors (note that colimit-preserving functors are exact, by [Lur, Prop. 1.1.4.1]).
On the spectral side, we have the stack of $L$-parameters $Z^{1}(W_{E},\widehat{G})/\widehat{G}$ over $\overline{\mathbb{Q}}_{\ell}$, defined in [FS21, §VIII]. Since it will figure extensively in this paper, we simplify the notation by defining
@@
[…]
@@
called the spectral action.
@@
$\operatorname{Rep}((\widehat{G}\rtimes Q)^{I})\to\operatorname{Perf}(\mathfrak{X})^{BW_{E}^{I}}\to\operatorname{End}_{\overline{\mathbb{Q}}_{\ell}}(\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})^{\omega})^{BW_{E}^{I}},$
@@
to the geometric Bernstein center $\mathcal{Z}^{\mathrm{geom}}(G)$ of $G$, landing inside the subring $\mathcal{Z}^{\mathrm{geom}}_{\mathrm{Hecke}}(G)$ of endomorphisms equivariant for the Hecke action [FS21, Def. IX.0.2]. For any object $A\in\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ we have a ring homomorphism $F_{A}:\mathcal{Z}^{\mathrm{spec}}(G)\to\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)$. When $A$ is Schur (i.e. $\operatorname{End}_{D_{\mathrm{lis}}(\mathrm{Bun}_{G})}(A)=\overline{\mathbb{Q}}_{\ell}$), the kernel of $F_{A}$ corresponds to a $\overline{\mathbb{Q}}_{\ell}$-point of $X$, which we refer to as the Fargues–Scholze parameter of $A$. By [Zou22, Theorem 5.2.1], this construction is compatible with that given in [FS21, §IX.4]. For $b\in B(G)$ and $\rho$ an irreducible admissible representation of $G_{b}(E)$, we will write “the Fargues–Scholze parameter of $\rho$” to mean the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ [^2] We note that using $i_{\natural}^{b}\rho$ instead of $i^{b}_{\ast}\rho$, as is done in [FS21, IX.7], produces the same result (e.g. by the argument in the proof of Proposition 2.7). Moreover, the Fargues–Scholze parameter of $i^{b}_{\ast}\rho$ is the composite of the “correct” Fargues–Scholze parameter of $\rho$ (defined using $i^{1}_{!}\rho$ in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G_{b}})$ as in [FS21, Def. IX.7.1]) and the twisted $L$-embedding $\widehat{G}_{b}\rtimes Q\to\widehat{G}\rtimes Q$, by [FS21, Thm IX.7.2]..
Now consider a derived stack $\mathfrak{Y}$ with a map $f:\mathfrak{Y}\to\mathfrak{X}$, which induces a symmetric monoidal functor $f^{\ast}:\operatorname{QCoh}(\mathfrak{X})\to\operatorname{QCoh}(\mathfrak{Y})$. Using the spectral action and $f^{\ast}$, we can then form the tensor product
$\operatorname{QCoh}(\mathfrak{Y})\otimes_{\operatorname{QCoh}(\mathfrak{X})}\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$
@@
To keep the definitions of this paper close to its theorems, we make our definitions in a restricted setting, where things are more concrete. Consider the character variety $X$ from above. Let $Y$ be a closed subscheme of $X$, and consider its derived pullback $\mathfrak{Y}$ to $\mathfrak{X}$. The structure sheaf $\mathcal{O}_{\mathfrak{Y}}$ is a commutative algebra object of $\operatorname{QCoh}(\mathfrak{X})$. Similarly, the structure sheaf $\mathcal{O}_{Y}$ of $Y$ is a commutative algebra object of $\operatorname{QCoh}(X)$, and pullback gives a symmetric monoidal functor $\operatorname{QCoh}(X)\to\operatorname{QCoh}(\mathfrak{X})$ sending $\mathcal{O}_{Y}$ to $\mathcal{O}_{\mathfrak{Y}}$. In particular, we may and will also view $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ as module for $\operatorname{QCoh}(X)$.
@@
Remark 2.4.
We note that our $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$ differs from the category $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})_{\varphi}$ defined by the first author in [HL24, Definition A.1]. An analogy is the following: If $A$ is a commutative ring and $\mathfrak{m}\subseteq A$ is a maximal ideal, then $\mathcal{D}(A/\mathfrak{m})$ is to $\mathcal{D}(A)$ as our $\mathcal{D}_{\mathrm{lis}}^{\varphi}(\mathrm{Bun}_{G})$ is to $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$, whereas $\mathcal{D}(\widehat{A}_{\mathfrak{m}})$ is to $\mathcal{D}(A)$ as $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})_{\varphi}$ is to $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$. Here $\widehat{A}_{\mathfrak{m}}$ is the complete local ring of $A$ at $\mathfrak{m}$, and $\mathcal{D}(\dots)$ denotes the unbounded derived category of a ring (viewed as a stable $\infty$-category). We will not use the categories $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})_{\varphi}$ in this paper.
@@
We start with part (1). For any $k$-algebras $R$ and $S$, the $\infty$-category of colimit-preserving functors $\mathcal{D}(R)\to\mathcal{D}(S)$ is equivalent to $\mathcal{D}(S\otimes_{k}R^{op})$, or equivalently to the derived $\infty$-category $\mathcal{D}(S,R)$ of $(S,R)$-bimodules for which the $k$-module structures coming from $S$ and $R$ agree. Here $X\in\mathcal{D}(S,R)$ corresponds to the functor $M\mapsto X\otimes_{R}^{L}M$, and for a given $F:\mathcal{D}(R)\to\mathcal{D}(S)$ the corresponding $X$ is $F(R)$, with the left $R^{op}$-module structure coming from the map
@@
Since $B$ is concentrated in degree 0, $\operatorname{End}_{\mathcal{D}(B,B)}(B)$ is coconnective. Since $A$ is also concentrated in degree 0, this means that the map $A\to\operatorname{End}_{\mathcal{D}(B,B)}(B)$ factors through the center $Z(B)=\pi_{0}(\operatorname{End}_{\mathcal{D}(B,B)}(B))$, giving the desired map $f$. Translating back from $\mathcal{D}(B,B)$ to $\operatorname{End}_{k}^{L}(\mathcal{D}(B))$ then gives the desired formula.
For part (2), by shifting we may assume that $M$ has homology in degree 0, and so lies in the heart of $\mathcal{D}(B)$ with respect to the usual t-structure. The assumption then says that $M$ has a canonical $B/IB$-module structure in the usual sense. Restriction along the canonical map $A/I\otimes^{L}_{A}B\to B/IB$ then gives a canonical $A/I\otimes_{A}^{L}B$-module structure, as desired. ∎